# line of sight CSI channel state information

### v2

@@
Upper Mid-Band Spectrum for 6G: Vision, Opportunity and Challenges
Ahmad Bazzi, Roberto Bomfin, Marco Mezzavilla, Sundeep Rangan, Theodore Rappaport, Marwa Chafii Ahmad Bazzi, Roberto Bomfin, and Marwa Chafii are with the Engineering Division, New York University Abu Dhabi (NYUAD), 129188, UAE (email: {ahmad.bazzi,roberto.bomfin, marwa.chafii}@nyu.edu)Ahmad Bazzi, Marwa Chafii, Sundeep Rangan, and Theodore Rappaport are with NYU WIRELESS, NYU Tandon School of Engineering, Brooklyn, 11201, NY, USA (email: {srangan,tsr}@nyu.edu).Marco Mezzavilla is with the Telecommunications Engineering at the Department of Electronics, Information, and Bioengineering (DEIB) at Politecnico di Milano, Milan, Italy. (email: marco.mezzavilla@polimi.it).
@@ Abstract
Driven by the pursuit of gigabit-per-second data speeds for future 6G mobile networks, in addition to the support of sensing and artificial intelligence applications, the industry is expanding beyond crowded sub-6 GHz bands with innovative new spectrum allocations. In this paper, we chart a compelling vision for 6G within the frequency range 3 (FR3) spectrum, i.e. $7.125$-$24.25$ $\operatorname{GHz}$, by delving into its key enablers and addressing the multifaceted challenges that lie ahead for these new frequency bands. Here we highlight the physical properties of this never-before used spectrum by reviewing recent channel measurements for outdoor and indoor environments, including path loss, delay and angular spreads, and material penetration loss, all which offer insights that underpin future 5G/6G wireless communication designs. Building on the fundamental knowledge of the channel properties, we explore FR3 spectrum agility strategies that balance coverage and capacity (e.g. data rate) tradeoffs, while also examining coexistence with incumbent systems, such as satellites, radio astronomy, and earth exploration. Moreover, we discuss the potential of massive multiple-input multiple-output, compact and digital architectures, and evaluate the potential of multiband sensing for FR3 integrated sensing and communications. Finally, we outline 6G standardization features that are likely to emerge from 3GPP radio frame innovations and open radio access network developments.
@@
$6$G, upper midband, FR3, channel characteristics, spectrum agility, ISAC
@@ I Introduction
By $2034$, global mobile data traffic is expected to grow by $5$x–$9$x, with artificial intelligence (AI) accounting for one-third of the traffic. There is an industry-wide consensus that $6$G is expected to launch around $2030$ with new spectrum in frequency range 3 (FR3) [1], which spans $7.125$-$24.25$ $\operatorname{GHz}$ [2]. To meet the $6$G timeline, it is crucial to allocate the necessary spectrum a few years in advance, in order to ensure incumbents are accommodated and that the technology is ready for mass adoption at the global launch of $6$G. Companies are actively conducting proof-of-concept product trials to address bottlenecks and tackle myriad implementation challenges. At the same time, regulators must determine how the spectrum will be allocated, and operators need to develop viable business cases and deployment/integration strategies for nationwide $6$G networks.
@@ I-A The FR3 Spectrum
As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3]. Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig. 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers.
The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S. National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services. Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
@@ II FR3 Channel Characteristics
Env Freq. (GHz) Dist (m) LOS NLOS LOS Omni RMS DS NLOS Omni RMS DS LOS Omni RMS ASA NLOS Omni RMS ASA
Omni PLE Sigma Omni PLE Sigma
InH [1] 6.75 11–97 1.34 3.51 2.72 9.21 37.7 48 40.9 58.2
16.95 11–97 1.32 2.66 3.05 8.11 22.1 40.7 34.2 43.5
28 4–46 1.2 [6] 1.8 2.7 [6] 9.7 10.8 17.1 Med: 39.1 Med: 31.8
UMi [4] 6.75 40–1000 1.79 2.57 2.56 6.53 62.8 75.6 16.79 25.62
16.95 40–1000 1.85 4.05 2.59 8.78 46.5 65.8 18.23 19.08
28 31–187 2.02 - 2.1 [7] 8.98 3.4-3.56[7] 8.91 26.7 46.3 Med:14 Med:30
InF [5] 6.75 9–38 1.39 1.86 1.78 2.46 14 30.3 23.71 66.38
16.95 9–38 1.75 3.1 2.11 3.29 12.7 29 14.32 50.4
@@ II-A FR3 Channel Modeling
The channel model introduced in TR $38.901$ as part of the 3GPP global standard body designed to encompass the entire frequency range from $0.5$ to $100$ $\operatorname{GHz}$ [8]. Integrated channel models are indispensable for addressing diverse propagation scenarios, ranging from urban environments to free space, yet 3GPP developed the TR 38.901 channel models without many field measurements across much of the $100\operatorname{GHz}$ wide swath of the spectrum, and without any measurements from the FR3 bands. As a result, the model is an estimate, formulated from sparse measurements across particular bands within the $0.5$ to $100\operatorname{GHz}$ range. Channel modeling research for spectra above $6\operatorname{GHz}$ was completed in Release 17 of TR $38.901$ in April $2022$, offering a comprehensive set of models for evaluating various physical layer technologies. However, the primary focus of $5$G channel modeling has been on frequencies below $6\operatorname{GHz}$ (FR1) and above $24\operatorname{GHz}$ (FR3). To address this gap, frequency interpolation techniques were used by 3GPP to estimate channel parameters in the $7$-$24\operatorname{GHz}$ band. To establish a more accurate channel model for the FR3 band, it is crucial to validate the TR 38.901 model and carefully examine the details of FR3 channel parameterization. For this, true empirical characterization of the many channel parameters is required. 3GPP channel parameters include pathloss exponents, penetration losses, shadowing effects, cross-polarization discrimination (also referred to as cross-polarization isolation), in addition to delay and angular spreads across different environments, such as indoor hotspot (InH), indoor factory (InF) and urban microcell (UMi) under various conditions, such as line of sight (LoS) and non line of sight (NLoS) with minimal outages.
@@
The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g. the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g. $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.95\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1]. In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE $=2$) and with less loss at lower frequencies. Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [9], [10]. NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g. NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.95\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [5], and $2.11$ at $16.95\operatorname{GHz}$ [5], the loss over distance in NLoS locations is lower (e.g., $12$ to $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g. NLoS PLE of $3.4$ to $3.56$ at $28\operatorname{GHz}$ UMi [7], highlighting how FR3 offers improved coverage in NLOS channels relative to higher-frequency mmWave bands. It is important to note, however, that practical systems use higher antenna gains at higher frequencies, and a common error is to attempt to use equivalent isotropic radiated power (EIRP) to compare coverage at different frequencies, since the transmit and receive antenna physical apertures are proportionally reduced with shrinking wavelength if antenna gains are left constant over frequency, such that equal power transmitters with EIRP over different frequency will give the false impression that higher bands have far less coverage. In practice, the proper comparison is to consider equal antenna aperture areas at different frequencies, where in practice the extra array gain at mmWave can overcome the additional channel loss by many dB [11].
@@
Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
@@
The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods. Measurements in [1], [4], [5] show that root mean square (RMS) delay spread (DS) decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g. $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF. The range of RMS DS was found to be $12.7-37.7\operatorname{ns}$ across InF, UMi and InH. DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate position location and timing methods in dense environments. Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e. $16.95\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming. Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
When considering the design and comparison of air interface standards, such as the 3GPP TR 38.901 statistical channel model, the channel measurements [4, 5] conducted at new FR3 frequencies suggest that no major changes are needed in terms of numerology compared to existing 5G- new radio (NR) protocols; however, the beam management may need modifications due to the different temporal and spatial statistics and the different number of antenna elements at both gNodeB (gNB) and user equipment (UE). In particular, directional links rely on precise angular alignment of transmitter and receiver beams, a process facilitated through beam management operations. For next-generation cellular networks, mechanisms are being discussed in 3GPP technical documents that address beam management, such as R1111-$2409673$ in November $18-22$ as part of 3GPP technical specification group (TSG) radio access network (RAN) working group (WG1), in addition to other 3GPP contributions such as R1111-$2409710$, R1111-$2409741$ and more. Directional combining in beam management dramatically enhance link gains and coverage distances, as first shown in [12], thereby maintaining acceptable communication quality when beam combining is available – something that both FR2 and FR3 will support. However, as shown in [4] , as the carrier frequency increases, the angular spread tends to decrease. In addition, the empirical RMS angular spread of arrival (ASA) at both $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$ was noticeably lower than the 3GPP model predictions [4]. Factors such as mobility, device rotation, and wide angular spread channels can impact beam alignment. In dynamic scenarios, beam measurements can quickly become outdated, leading to inaccurate beam selection and performance degradation, which further impacts signal-to-interference-plus-noise ratio (SINR) and spectral efficiency (SE). The newly reported angular spreads in [4] implies that modifications are required for beam management in the new goldilocks FR3 mid-band spectrum. For sensing applications in ISAC, this also means that precise and accurate beam alignment is needed for target detection, localization and tracking. In addition to channel modeling, the call for new metrics and figures-of-merit, such as the Waste Factor, offer potential for energy efficiency optimization in $6$G wireless systems.
II-B FR3 frequency-dependent features and losses

### v3

@@
IEEE
Institute of Electrical and Electronics Engineers
IMT
International Mobile Telecommunication
CRLB
Cramér-Rao lower bound
RSS
received signal strength
DMC
dense multipath components
$E$
energy efficiency
Upper Mid-Band Spectrum for 6G:
Vision, Opportunity and Challenges
Ahmad Bazzi, Roberto Bomfin, Marco Mezzavilla, Sundeep Rangan, Theodore S. Rappaport, Marwa Chafii Ahmad Bazzi, Roberto Bomfin, and Marwa Chafii are with the Engineering Division, New York University Abu Dhabi (NYUAD), 129188, UAE (email: {ahmad.bazzi,roberto.bomfin, marwa.chafii}@nyu.edu)Ahmad Bazzi, Marwa Chafii, Sundeep Rangan, and Theodore Rappaport are with NYU WIRELESS, NYU Tandon School of Engineering, Brooklyn, 11201, NY, USA (email: {srangan,tsr}@nyu.edu).Marco Mezzavilla is with the Telecommunications Engineering at the Department of Electronics, Information, and Bioengineering (DEIB) at Politecnico di Milano, Milan, Italy. (email: marco.mezzavilla@polimi.it).
@@ Abstract
Driven by the pursuit of gigabit-per-second data speeds for future 6G mobile networks, in addition to the support of sensing and artificial intelligence applications, the industry is expanding beyond crowded sub-6 GHz bands with innovative new spectrum allocations. In this paper, we chart a compelling vision for 6G within the frequency range 3 (FR3) spectrum, i.e. 7.1257.125-$24.25$ $\operatorname{GHz}$, by delving into its key enablers and addressing the multifaceted challenges that lie ahead for these new frequency bands. Here we highlight the physical properties of this never-before used spectrum for cellular by reviewing recent channel measurements for outdoor and indoor environments, including path loss, delay and angular spreads, and material penetration loss, all which offer insights that underpin future 5G/6G wireless communication designs. Building on the fundamental knowledge of the channel properties, we explore FR3 spectrum agility strategies that balance coverage and capacity tradeoffs, while examining coexistence with incumbent systems, such as satellites, radio astronomy, and earth exploration. Moreover, we discuss the potential of massive multiple-input multiple-output technologies, challenges for commercial deployment, and potential solutions for FR3, including multiband sensing for FR3 integrated sensing and communications. Finally, we outline 6G standardization features that are likely to emerge from 3GPP radio frame innovations and open radio access network developments.
@@
66G, upper midband, FR3, channel characteristics, spectrum agility, ISAC
@@ I Introduction
By $2034$, global mobile data traffic is expected to grow by five- to nine-fold, with artificial intelligence (AI) accounting for one-third of the traffic. There is an industry-wide consensus that 66G is expected to launch around $2030$ with new spectrum in frequency range 3 (FR3) [1], which spans 7.1257.125-$24.25$ $\operatorname{GHz}$ [2]. To meet the 66G timeline, it is crucial to allocate the necessary spectrum a few years in advance, in order to ensure incumbents are accommodated and that the technology is ready for mass adoption at the global launch of 66G. Companies are actively conducting proof-of-concept product trials to address bottlenecks and tackle myriad implementation challenges. At the same time, regulators must determine how the spectrum will be allocated, and operators need to develop viable business cases and deployment/integration strategies for nationwide 66G networks.
@@ I-A The FR3 Spectrum
As the global wireless industry advances towards 66G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3]. Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for 66G, FR3 frequencies (7.1257.125-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig. 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers. U.S. regulators have signalled that 7.257.25-$7.4$ GHz will serve as an initial slice for FR3, which is favorable as it avoids the $22$ GHz water-vapour and $60$ GHz oxygen absorption peaks, keeping rain attenuation modest enough for macro-cell coverage. Further FR3 spectrum is expected to be layered outward from this starter band.
The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S. National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services. Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
@@
I-B Improvement items by ITU, 3GPP & Organization
3GPP work started in 2024 during Release 19, including channel models for the FR3. The recent FR3 measurements by NYU WIRELESS are unified in Section II. Meanwhile, the ITU has identified a list of capabilities in International Mobile Telecommunication (IMT)-$2030$ in the form of improvements and additional capabilities compared to those in IMT-$2020$. The improvements include: a peak data rate of up to $200\operatorname{Gbps}$, which is $10\times$ that of IMT-$2020$, in addition to user experienced data rates, which are $300\operatorname{Mbps}$ and $500\operatorname{Mbps}$, i.e a factor of $3$ to $5$ improvement. The area traffic capacity are expected to be $30$ and $50\operatorname{Mbps}$ per square meter and the spectrum efficiency was set to an improvement of $1.5$ and $3$ higher than that of IMT-$2020$, while connection density is set to a target of $10^{6}$-$10^{8}$ devices per km2. In Section III, we discuss different methods to achieve the aforementioned factors. Moreover, we highlight related challenges for commercial deployment in Section IV. Moreover, IMT-2030 also calls for enhancing radio access network (RAN), including flexibility, intelligence, and resiliency, so that a single RAN can be sliced, AI-optimized, and seamlessly span terrestrial and non-terrestrial links while guaranteeing specific QoS. Besides RAN, we mention several open questions and provide potential solutions in Section V.
@@ II-A FR3 Channel Modeling
The channel model introduced in TR $38.901$ as part of the 3GPP global standard body designed to encompass the entire frequency range from $0.5$ to $100$ $\operatorname{GHz}$ [6]. Integrated channel models are indispensable for addressing diverse propagation scenarios, ranging from urban environments to free space, yet 3GPP developed the TR 38.901 channel models without many field measurements across much of the $100\operatorname{GHz}$ wide swath of the spectrum, and without any measurements from the FR3 bands. Consequently, the model is an estimate, formulated from sparse measurements across particular bands within the $0.5$ to $100\operatorname{GHz}$ range.
@@
Channel modeling research for spectra above $6\operatorname{GHz}$ was completed in Release 17 of TR $38.901$ in April $2022$, offering a comprehensive set of models for evaluating various physical layer technologies. However, the primary focus of 55G channel modeling has been on frequencies below $6\operatorname{GHz}$ (FR1) and above $24\operatorname{GHz}$ (FR3). To address this gap, frequency interpolation techniques were used by 3GPP to estimate channel parameters in the $7$-$24\operatorname{GHz}$ band. To establish a more accurate channel model for the FR3 band, it is crucial to validate the TR 38.901 model and carefully examine the details of FR3 channel parameterization. For this, true empirical characterization of the many channel parameters is required.
II-B FR3 pathloss exponents over different environments
@@
The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional PLE values, synthesized from directional channel measurements, in line of sight (LoS) scenarios were slightly lower than those in millimeter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g. the omnidirectional LoS PLEs in indoor hotspot (InH), urban microcell (UMi) and indoor factory (InF) for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g. $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.95\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1]. In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE is $2$) and with less loss at lower frequencies. Although antenna patterns and gains vary widely, and higher frequencies enable usage of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [8], [9]. Omnidirectional non line of sight (NLoS) PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g. NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.95\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [5], and $2.11$ at $16.95\operatorname{GHz}$ [5], the loss over distance in NLoS locations is lower (e.g., $12$ to $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g. NLoS PLE of $3.4$ to $3.56$ at $28\operatorname{GHz}$ UMi [7], highlighting how FR3 offers improved coverage in NLOS channels relative to higher-frequency mmWave bands. It is important to note, however, that practical systems use higher antenna gains at higher frequencies, and a common error is to attempt to use equivalent isotropic radiated power (EIRP) to compare coverage at different frequencies, since the transmit and receive antenna physical apertures are proportionally reduced with shrinking wavelength if antenna gains are left constant over frequency, such that equal power transmitters with EIRP over different frequency will give the false impression that higher bands have far less coverage. In practice, the proper comparison is to consider equal antenna aperture areas at different frequencies, where in practice the extra array gain at mmWave can overcome the additional channel loss by many dB [10]. In this work, Fig. 2 and Fig. 3 do not consider the reality of greater channel gain at higher frequencies, and hence overestimate the achieavable capacity of lower frequencies and underpredict higher frequency rates.
@@
Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying 66G technologies.
II-C FR3 delay spreads, angular spreads & penetration losses
@@
The DS of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods. Measurements in [1], [4], [5] show that root mean square (RMS) DS decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g. $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF. The range of RMS DS was found to be $12.7$-$37.7\operatorname{ns}$ across InF, UMi and InH. DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate location and timing methods in dense environments.
Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e. $16.95\operatorname{GHz}$, exhibited a narrower RMS AS compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming.
Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
II-D FR3 frequency-dependent features and losses